OpenStudy (anonymous):

Simplify the rational expression. State any restrictions on the variable. t^2-4t-32/t-8 a. -t - 4; t = 8 b. -t + 4; t = -8 c. t + 4; t = 8

OpenStudy (tyteen4a03):

Hint: \(t^2-4t-32 = (t-8)(t+4)\)

OpenStudy (anonymous):

b?

OpenStudy (tyteen4a03):

No, it's c. After simplifying, you should get t+4.

OpenStudy (jiteshmeghwal9):

factor the numerator t^2-4t-32=t^2-8t+4t-32 =t(t-8)+4(t-8) =(t+4)(t-8)

OpenStudy (jiteshmeghwal9):

now put the value instead of numerator

OpenStudy (jiteshmeghwal9):